# SDEs with supercritical distributional drifts

### v1

@@ Abstract.
This work is partially supported by NNSFC grants of China (Nos. 12131019), and the German Research Foundation (DFG) through the Collaborative Research Centre(CRC) 1283/2 2021 - 317210226 “Taming uncertainty and profiting from randomness and low regularity in analysis, stochastics and their applications”.
@@ 1. Introduction
Throughout this paper we fix $d\geqslant 2$. Let $(W_{t})_{t\geqslant 0}$ be a standard $d$-dimensional Brownian motion on a stochastic basis $(\Omega,{\mathscr{F}},({\mathscr{F}}_{t})_{t\geqslant 0},{\mathbf{P}})$. Consider the following stochastic differential equation (SDE) in ${\mathbb{R}}^{d}$ driven by Brownian motion $W$:
${\mathord{{\rm d}}}X_{t}=b(t,X_{t}){\mathord{{\rm d}}}t+\sqrt{2}{\mathord{{\rm d}}}W_{t},\ \ t\geqslant 0,$ (1.1)
where the drift $b$ is a time-dependent distribution. Since $b$ is a distribution, the drift term does not make a meaning in the classical sense as we cannot assign a value to a distribution at the point $X_{t}$. To formulate a solution, one natural approach is to employ mollifying approximation. Let $\phi_{n}(x)=n^{d}\phi(nx)$ be a family of modifiers, where $\phi\in C^{\infty}_{c}({\mathbb{R}}^{d})$ is a smooth probability density function with compact support. Define the smooth approximation of $b$ by convolution as:
@@
We begin by introducing the concept of weak solutions for the above SDE with a distributional drift. Let ${\mathcal{P}}({\mathbb{R}}^{d})$ be the space of all probability measures on ${\mathbb{R}}^{d}$.
@@
Let $\mathfrak{F}:=(\Omega,{\mathscr{F}},({\mathscr{F}}_{t})_{t\geqslant 0},{\mathbf{P}})$ be a stochastic basis and $(X,W)$ be a pair of ${\mathbb{R}}^{d}$-valued continuous ${\mathscr{F}}_{t}$-adapted processes on $\mathfrak{F}$. We call $(\mathfrak{F},X,W)$ a weak solution of SDE (1.1) with initial distribution $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$ if $W$ is an ${\mathscr{F}}_{t}$-Brownian motion, ${\mathbf{P}}\circ X_{0}^{-1}=\mu$, and for all $t\in[0,T]$,
$X_{t}=X_{0}+A^{b}_{t}+\sqrt{2}W_{t},\ \ a.s.,$
where $A^{b}_{t}:=\lim_{n\to\infty}\int^{t}_{0}b_{n}(s,X_{s}){\mathord{{\rm d}}}s$ exists in the $L^{2}$-sense.
It is noted that the definition of a weak solution in Definition 1.1 depends on the choice of mollifiers $\phi_{n}$. The goal of this paper is to find conditions on $b$ as weak as possible so that SDE (1.1) can be solved in the sense of Definition 1.1 and to construct the associated Markov process.
To present our results, we commence by conducting a basic scale analysis, which enables us to distinguish between subcritical, critical, and supercritical drifts. Let ${\mathbf{H}}^{\alpha}_{p}$ (resp. $\dot{\mathbf{H}}^{\alpha}_{p}$) be the (resp. homogenous) Bessel potential space, where $\alpha\in{\mathbb{R}}$ and $p\in[1,\infty]$ (see [5] or Subsection 2.1 below for precise definitions). Suppose that for some $\alpha\in{\mathbb{R}}$ and $q,p\in[1,\infty]$,
$b\in L^{q}({\mathbb{R}}_{+};\dot{\mathbf{H}}^{\alpha}_{p}),$
and SDE (1.1) admits a solution denoted by $X$. For $\lambda>0$, define
$X^{\lambda}_{t}:=\lambda^{-1}X_{\lambda^{2}t},\ \ W^{\lambda}_{t}:=\lambda^{-1}W_{\lambda^{2}t},\ \ b_{\lambda}(t,x):=\lambda b(\lambda^{2}t,\lambda x).$
Formally, one sees that
${\mathord{{\rm d}}}X^{\lambda}_{t}=b_{\lambda}(t,X^{\lambda}_{t}){\mathord{{\rm d}}}t+\sqrt{2}{\mathord{{\rm d}}}W^{\lambda}_{t}.$
Moreover, by the change of variable we have
@@
As $\lambda\to 0$, we categorize SDE (1.1) into the following three cases:
$\mbox{{\bf Subcritical}: $\frac{d}{p}+\frac{2}{q}<1+\alpha$};\ \mbox{{\bf Critical}: $\frac{d}{p}+\frac{2}{q}=1+\alpha$};\ \mbox{{\bf Supercritical}: $\frac{d}{p}+\frac{2}{q}>1+\alpha$}.$
@@
Let’s perform a straightforward analysis of the differentiability index $\alpha$ when $b\in{\mathbf{C}}^{\alpha}$ with $\alpha<0$. According to the Schauder theory of the heat equation, $u$ belongs to at most $\mathbf{C}^{2+\alpha}$. To make the product $b\cdot\nabla u$ meaningful, we need to stipulate that $01+2\alpha>$ 0, which implies $\alpha>-\frac{1}{2}$.
We summarize part of well-known results regarding the SDE (1.1) in the three cases, categorized based on the value of $\alpha$. For clarity, we will utilize the following abbreviations in the subsequent table:
@@
    When $\alpha=0$ and $p=q=\infty$, Veretenikov [41] firstly obtained the strong well-posedness of SDE (1.1) for any starting point $X_{0}$.
@@
    In the subcritical case with $\alpha=0$, Krylov and Röckner [29] demonstrated the strong well-posedness of SDE (1.1) using Girsanov’s technique. Related results obtained through Zvonkin’s transformation can be found in [43, 44].
@@
    In the critical case with $\alpha=0$, the weak and the strong well-posedness are studied in recent work [4, 24, 33, 34]. See also Krylov’s series of works [26, 27, 28].
@@
    In the supercritical case, when $\alpha=0$ and $\mathord{{\rm div}}b=0$, a weak solution is constructed in [49] using the maximum principle proven by De-Giorgi’s iteration technique method. For the case of multiplicative and possibly degenerate noise, see [46].
@@
    When $\alpha\in(-\frac{1}{2},0)$ and $b$ is a time-independent vector field belonging to the space $\mathbf{C}^{\alpha}$, the existence and uniqueness of a solution termed as “virtual solution” was proved in [14]. When $\alpha\in[-\frac{1}{2},0)$ and $b\in{\mathbf{H}}^{\alpha}_{p}$ with $\alpha\in[-\frac{1}{2},0]$ and $p>\frac{d}{1+\alpha}$, the authors in [48] showed that there is a unique weak solution $(\mathfrak{F},X,W)$ to SDE (1.1) in the class that the following Krylov estimate holds: for some $\theta,T>0$ and any $m\in{\mathbb{N}}$, $f\in C_{c}({\mathbb{R}}^{d})$ and $0\leqslant t_{0}<t_{1}\leqslant T$,
    $\left\|\int^{t_{1}}_{t_{0}}f(X_{s}){\mathord{{\rm d}}}s\right\|_{L^{m}(\Omega)}\leqslant C(t_{1}-t_{0})^{\frac{1+\theta}{2}}\|f\|_{{\mathbf{H}}^{\alpha}_{p}},$ (1.4)
    where the constant $C=C(m,T,d,\alpha,p,\|b\|_{{\mathbf{H}}^{\alpha}_{p}})>0$. In particular, the above Krylov estimate implies that $A^{b}_{t}:=\lim_{n\to\infty}\int^{t}_{0}b_{n}(X_{s}){\mathord{{\rm d}}}s$ exists and $A^{b}_{\cdot}$ is a zero energy process. In the one-dimensional case, the zero energy solution was explored by Bass and Chen in [2].
@@
    In the subcritical case, with $\alpha\in(-1,0]$ and a time-independent, divergence-free $b$, a unique energy solution on the torus is established in [20] for SDE (1.1) with initial data whose probability distribution has an $L^{2}$-density with respect to the Lebesgue measure.
@@
    In the critical and supercritical cases, when $b=b(x)\in\mathbf{H}^{-1}_{p}$ with some $p>2$ and is divergence-free, the uniqueness of the energy solution is established in [19] for any initial data whose law has a bounded density with respect to the Lebesgue measure.
  • $\bullet$
    Without assuming that $b$ is divergence-free, when $b\in\mathbf{C}^{\alpha}$ for some $\alpha\in(-\frac{2}{3},-\frac{1}{2})$, representing a certain Gaussian noise, the authors in [11] and [7] utilized the rough path and paracontrolled theory to independently establish the existence of a unique martingale solution. Additional references to this scenario can be found in [25] for the Lévy process case and [22] for the degenerate kinetic case.
@@ 1.1. Main results
In this subsection, we present our main results in two cases: the subcritical case and the supercritical case. Our first result establishes the well-posedness of SDE (1.1) in the subcritical case when $b$ is divergence-free.
@@
Let $\alpha_{b}\in(-1,-\frac{1}{2}]$ and $p_{b},q_{b}\in[2,\infty]$ satisfy $\frac{d}{p_{b}}+\frac{2}{q_{b}}<1+\alpha_{b}$. Suppose that $b\in\cap_{T>0}{\mathbb{L}}^{q_{b}}_{T}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}$ is divergence-free. Then for any $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$, there is a weak solution to SDE (1.1) starting from initial distribution $\mu$ in the sense of Definition 1.1, which is also unique in the class that the following Krylov estimate holds: for any $(\alpha,p,q)\in[\alpha_{b},0]\times[2,\infty]^{2}$ with
$\tfrac{q_{b}}{2}\leqslant q\leqslant q_{b},\ \ p\leqslant p_{b},\ \ \alpha-\tfrac{d}{p}-\tfrac{2}{q}\geqslant\alpha_{b}-\tfrac{d}{p_{b}}-\tfrac{2}{q_{b}},$ (1.5)
for any $T>0$, there is a $\theta>0$ such that for any $m\in{\mathbb{N}}$, $f\in{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}\cap{\mathbb{L}}^{q}_{T}C^{\infty}_{b}$ and $0\leqslant t_{0}<t_{1}\leqslant T$,
$\left\|\int^{t_{1}}_{t_{0}}f(s,X_{s}){\mathord{{\rm d}}}s\right\|_{L^{m}(\Omega)}\lesssim_{C}(t_{1}-t_{0})^{\frac{1+\theta}{2}}\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}},$ (1.6)
where the constant $C$ only depends on $\theta,m,T,d,\alpha,p,q,\alpha_{b},p_{b},q_{b}$ and $\|b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}}$. Moreover, for each $t>0$ and $X_{0}=x\in{\mathbb{R}}^{d}$, the law of $X_{t}$ admits a density $\rho_{t}(x,x^{\prime})$ called the heat kernel of $\Delta+b\cdot\nabla$, which enjoys the following two sides Gaussian estimate: for fixed $T>0$, all $x,x^{\prime}\in{\mathbb{R}}^{d}$ and $t\in(0,T]$,
$C_{0}t^{-d/2}\mathrm{e}^{-c_{0}|x-x^{\prime}|^{2}/t}\leqslant\rho_{t}(x,x^{\prime})\leqslant C_{1}t^{-d/2}\mathrm{e}^{-c_{1}|x-x^{\prime}|^{2}/t},$
where $C_{0},C_{1},c_{0},c_{1}>0$ only depend on the parameters $T,d,\alpha_{b},p_{b},q_{b}$ and $\|b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}}$.
@@
In the subcritical case, our result improves upon [20, Theorem 2.10] since our initial distribution is not required to have an $L^{2}$-density; it can be a Dirac measure. Moreover, we are working in the whole space, not in the torus. Our proof is based on the Schauder estimate of heat equation in Besov spaces and Zvonkin’s transformation. Moreover, by (1.6), one easily sees that the solution does not depend on the choice of the mollifiers.
@@
(ii) ${\mathbf{B}}^{-1}_{\infty,2}\subset{\mathcal{B}}$. We shall prove it after introducing Besov spaces (see Lemma 2.7 below).
In the following we fix a time level $T>0$. Let ${\mathbb{C}}_{T}:=C([0,T];{\mathbb{R}}^{d})$ be the Banach space of all continuous functions which is endowed with the natural filtration ${\mathscr{B}}_{t}:=\sigma(\omega_{s},s\leqslant t)$. A path in ${\mathbb{C}}_{T}$ is denoted by $\omega$. The canonical process is denoted by
$w_{t}(\omega)=\omega(t).$
@@
Let $\alpha_{b}\in[-1,0]$ and $p_{b},q_{b}\in[2,\infty]$ satisfy $\frac{2}{q_{b}}+\frac{d}{p_{b}}<2+\alpha_{b}$. Suppose that $b\in{\mathbb{L}}^{q_{b}}_{T}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}$ is divergence-free. For any initial distribution $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$ having $L^{2}$-density $\rho_{0}$, there is a weak solution $(X,W)$ to SDE (1.1) in the sense of Definition 1.1, which has a density $\rho_{t}\in{\mathbb{L}}^{\infty}_{T}L^{2}\cap{\mathbb{L}}^{2}_{T}{\mathbf{H}}^{1}_{2}$ that satisfies the Fokker-Planck equation in the distributional sense
$\partial_{t}\rho=\Delta\rho-\mathord{{\rm div}}(b\rho).$
Moreover, consider the approximation SDE
@@
    Let ${\mathbb{P}}_{\mu}$ be the law of the solution constructed in (A). The following Markov property holds:
@@
    If in addition that $b=b_{1}+b_{2}$, where $b_{1}\in{\mathbb{L}}^{\infty}_{T}{\mathcal{B}}$ and $b_{2}\in{\mathbb{L}}^{2}_{T}L^{2}$, then without subtracting subsequence, ${\mathbb{P}}_{n}$ weakly converges to ${\mathbb{P}}_{\mu}$ in ${\mathbb{C}}_{T}$ as $n\to\infty$.
@@
In the supercritical case, we are unable to demonstrate the uniqueness of weak solutions a prioril. However, the assertion (A) affirms that we can identify a subsequence such that, for any initial value, the corresponding approximation solution converges weakly to a solution. This enables us to select a Markov process. In the case where $b=b_{1}+b_{2}$, with $b_{1}\in{\mathbb{L}}^{\infty}_{T}{\mathcal{B}}$ and $b_{2}\in{\mathbb{L}}^{2}_{T}L^{2}$, it is unique in the approximation sense since it is not necessary to choose a subsequence.
Remark 1.6.
One of the motivations for exploring the supercritical case arises from the observation of the super-diffusion phenomenon, as discussed in [8, 9]. In this context, the drift term is defined as the Leray projection of the 2-dimensional white noise, specifically represented as:
${\mathbf{E}}\big{(}b(f)b(g)\big{)}=\int_{{\mathbb{R}}^{2}}\hat{f}(\xi)\left({\mathbb{I}}_{2\times 2}-\frac{\xi\otimes\xi}{|\xi|^{2}}\right)\hat{g}(-\xi){\mathord{{\rm d}}}\xi,$
where $\hat{f}$ is the Fourier transform of Schwartz function $f:{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$. Given the absence of well-posedness results for the related stochastic differential equations as presented in [9], modifications were introduced to the drift term by truncating its Fourier transform, i.e.
${\mathbf{E}}\big{(}b_{\varepsilon}(f)b_{\varepsilon}(g)\big{)}=\int_{{\mathbb{R}}^{2}}1_{\{|\xi|<1/\varepsilon\}}\hat{f}(\xi)\left({\mathbb{I}}_{2\times 2}-\frac{\xi\otimes\xi}{|\xi|^{2}}\right)\hat{g}(-\xi){\mathord{{\rm d}}}\xi,$
where $\varepsilon$ is taken $1$ in [9]. Note that estimate (6.14) and Lemma 5.3 below allows us to establish the tightness of solutions to the SDE with the modified drift term $b_{\varepsilon_{n}}/\sqrt{\ln(1/\varepsilon_{n})}$, where $\varepsilon_{n}\to 0$.
@@
When $f=0$ and $b=\mathord{{\rm div}}a$, where $a:\mathbb{R}^{d}\to\mathbb{R}^{d}\otimes\mathbb{R}^{d}$ is an antisymmetric matrix-valued measurable function, the Harnack inequality was established in [32] for $a\in L^{\infty}$ and in [35] for $a\in BMO$. In the case where $u_{0}=0$ and $b\in L^{p}$ with $\mathord{{\rm div}}b=0$, through De-Giorgi’s argument, the maximum estimate was obtained in [49] for $p>d/2$ and in [46, Corollary 1.5] for $p>(d-1)/2$. In this work, for drifts that are distribution-valued and under certain conditions on the divergence of $b$, we show the weak well-posedness of PDE (1.8) using the energy method. To establish the existence of a weak solution to SDE (1.1), we rely on two types of Krylov’s estimates. One is used to prove tightness (see Lemma 5.1), while the other is utilized to take limits (see Lemma 5.5).
To establish conclusions (A) and (C), we will make use of the notion of generalized martingale solutions. In cases involving distributional drift, it’s important to note that the weak solution is not generally equivalent to the generalized martingale solutions. However, in our proofs, both of these solutions are the limits of smooth approximation solutions. It’s worth emphasizing that in the supercritical case, even without uniqueness, we have the flexibility to select a Markov process.
@@
In [23, Proposition 1], the entropy solution of the above Liouville equation is directly constructed. However, the existence of a solution to the SDE (1.9) appears to be an open question. In the context of the two-dimensional vortex model, where

### v3

@@ Abstract.
This work is partially supported by National Key R&D program of China (No. 2023YFA1010103) and NNSFC grants of China (Nos. 12131019), and the German Research Foundation (DFG) through the Collaborative Research Centre(CRC) 1283/2 2021 - 317210226 “Taming uncertainty and profiting from randomness and low regularity in analysis, stochastics and their applications”.
@@ 1. Introduction
Throughout this paper, we fix $d\geqslant 2$ and consider the following stochastic differential equation (SDE) in $\mathbb{R}^{d}$, driven by a standard $d$-dimensional Brownian motion $W$:
$\mathrm{d}X_{t}=b(t,X_{t})\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}W_{t},\quad t\geqslant 0,$ (1.1)
where the drift $b$ is a time-dependent distribution. Since $b$ is a distribution, the drift term is not meaningful in the classical sense, as it is not possible to assign a value to a distribution at the point $X_{t}$. To define solutions, a natural approach is to use mollifying approximations. Let $\phi_{n}(x)=n^{d}\phi(nx)$ be a family of mollifiers, where $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$ is a smooth probability density function with compact support. The smooth approximation of $b$ is then defined by convolution as follows:
@@
Let $\mathcal{P}(\mathbb{R}^{d})$ denote the space of all probability measures on $\mathbb{R}^{d}$. We begin by introducing the concept of weak solutions for the SDE (1.1) with a distributional drift.
@@
[…]
where $A^{b}_{t}:=\lim_{n\to\infty}\int_{0}^{t}b_{n}(s,X_{s})\,\mathrm{d}s$ exists in the $L^{2}$-sense.
To the best of the authors’ knowledge, the concept of weak solutions introduced above first appeared in [2] (see also [50]). It is important to note that the definition of a weak solution in Definition 1.1 depends on the choice of mollifiers $\phi_{n}$. The primary objective of this paper is to identify the weakest possible conditions on $b$ under which the SDE (1.1) admits a solution in the sense of Definition 1.1 and to construct the corresponding Markov process.
To present our results, we begin with a basic scaling analysis, which allows us to distinguish between subcritical, critical, and supercritical drifts. Let ${\mathbf{H}}^{\alpha}_{p}$ (resp. $\dot{{\mathbf{H}}}^{\alpha}_{p}$) denote the (resp. homogeneous) Bessel potential space, where $\alpha\in\mathbb{R}$ and $p\in[1,\infty]$ (see [5] or Subsection 2.1 below for precise definitions). Suppose that $b\in L^{q}(\mathbb{R}_{+};\dot{{\mathbf{H}}}^{\alpha}_{p})$ for some $\alpha\in\mathbb{R}$ and $q,p\in[1,\infty]$, and the SDE (1.1) admits a solution denoted by $X$. For $\lambda>0$, define
$X^{\lambda}_{t}:=\lambda^{-1}X_{\lambda^{2}t},\quad W^{\lambda}_{t}:=\lambda^{-1}W_{\lambda^{2}t},\quad b_{\lambda}(t,x):=\lambda b(\lambda^{2}t,\lambda x).$
Formally, it follows that
$\mathrm{d}X^{\lambda}_{t}=b_{\lambda}(t,X^{\lambda}_{t})\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}W^{\lambda}_{t}.$
Moreover, by a change of variables, we have
@@
As $\lambda\to 0$, we classify the SDE (1.1) into the following three cases:
[…]
@@
Let us perform a straightforward analysis of the differentiability index $\alpha$ when $b\in{\mathbf{C}}^{\alpha}$ with $\alpha<0$ (see Definition 2.2 below for a definition of ${\mathbf{C}}^{\alpha}$). According to the Schauder theory for the heat equation, $u$ belongs to at most $\mathbf{C}^{2+\alpha}$. To ensure the product $b\cdot\nabla u$ is well-defined, we need the condition $1+2\alpha>0$, which implies $\alpha>-1/2$.
We summarize some well-known results regarding the SDE (1.1) in three cases, categorized based on the value of $\alpha$. For clarity, we introduce the following abbreviations:
@@
    When $\alpha=0$ and $p=q=\infty$, Veretenikov [43] first established the strong well-posedness of the SDE (1.1) for any starting point $X_{0}$.
@@
    In the subcritical case with $\alpha=0$, Krylov and Röckner [31] demonstrated the strong well-posedness of the SDE (1.1) using Girsanov’s technique. Related results obtained through Zvonkin’s transformation are available in [45, 46].
@@
    In the critical case with $\alpha=0$, the weak and strong well-posedness were studied in recent works [4, 26, 35, 36]. See also Krylov’s series of works [29, 30].
@@
[…]
@@
    When $\alpha\in(-\frac{1}{2},0)$ and $b\in\mathbf{C}^{\alpha}$ is a time-independent vector field, the existence and uniqueness of a solution, termed a “virtual solution,” were proved in [13]. When $\alpha\in[-\frac{1}{2},0)$ and $b\in{\mathbf{H}}^{\alpha}_{p}$ with $\alpha\in[-\frac{1}{2},0]$ and $p>\frac{d}{1+\alpha}$, the authors in [50] showed that there is a unique weak solution $(\mathfrak{F},X,W)$ to the SDE (1.1) in the class such that the following Krylov estimate holds: for some $\theta,T>0$ and any $m\in\mathbb{N}$, $f\in C_{c}(\mathbb{R}^{d})$, and $0\leqslant t_{0}<t_{1}\leqslant T$,
    $\left\|\int_{t_{0}}^{t_{1}}f(X_{s})\,{\mathord{{\rm d}}}s\right\|_{L^{m}(\Omega)}\leqslant C(t_{1}-t_{0})^{\frac{1+\theta}{2}}\|f\|_{{\mathbf{H}}^{\alpha}_{p}},$ (1.4)
    where $C=C(m,T,d,\alpha,p,\|b\|_{{\mathbf{H}}^{\alpha}_{p}})>0$. In particular, the above Krylov estimate implies that $A^{b}_{t}:=\lim_{n\to\infty}\int_{0}^{t}b_{n}(X_{s})\,{\mathord{{\rm d}}}s$ exists, and $A^{b}_{\cdot}$ is a zero-energy process. In the one-dimensional case, the zero-energy solution was explored by Bass and Chen in [2].
@@
    In the subcritical case, with $\alpha\in(-1,0]$ and a time-independent, divergence-free $b$, the existence of a unique energy solution on the torus is established in [19] for the SDE (1.1), assuming initial data whose probability distribution has an $L^{2}$-density with respect to the Lebesgue measure. In the supercritical case, when $b=b_{1}+b_{2}$, where $b_{1}$ is divergence free and belongs to $L^{p}(\mathbb{R}_{+};{\mathbf{B}}^{-\gamma}_{p,1})$ for some $\gamma\in[0,1]$ and $p\geqslant 2/(1-\gamma)$, and $b_{2}$ represents some scaling-critical, non-divergence-free perturbations, the uniqueness of the energy solution is established in [20] for any initial data whose law is absolutely continuous with respect to the Lebesgue measure. See also [18] for related results.
@@
    Without assuming that $b$ is divergence-free, when $b\in\mathbf{C}^{\alpha}$ for some $\alpha\in\left(-\frac{2}{3},-\frac{1}{2}\right)$, the authors of [10] and [8] independently utilized rough path theory and paracontrolled distribution techniques to establish the existence of a unique martingale solution in the renormalized sense. Additional references include [27], which considers the case of Lévy processes, and [23], which addresses the degenerate kinetic case. It is important to note, however, that in their framework, not all distributions $b\in\mathbf{C}^{\alpha}$ with $\alpha\in\left(-\frac{2}{3},-\frac{1}{2}\right)$ can be treated. In fact, Theorem 6.7 in [24] provides a counterexample showing that, for any $\alpha\leqslant-\frac{1}{2}$, there exists a distribution $b\in\mathbf{C}^{\alpha}$ for which uniqueness fails under their definition of solutions.
@@ 1.1. Main results
In this subsection, we present our main results in two cases: the subcritical case and the supercritical case. Our first result establishes the well-posedness of SDE (1.1) in the subcritical case.
@@
Let $m\geqslant 2$, $\alpha_{b}\in(-1,-\frac{1}{2}]$, and $p_{b},q_{b}\in[2,\infty]$ satisfy
$\theta_{0}:=1+\alpha_{b}-\tfrac{d}{p_{b}}+\tfrac{2}{q_{b}}>0,\quad m\theta_{0}>2.$ (1.5)
Suppose that $b=b_{1}+b_{2}$, where $b_{1},\mathord{{\rm div}}b_{1}\in\cap_{T>0}{\mathbb{L}}^{q_{b}}_{T}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}$ and $|b_{2}(t,x)|\leqslant c_{0}+c_{1}|x|$ for some $c_{0},c_{1}\geqslant 0$.
  1. (i)
    For any $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$ with finite $m$-order moment, there exists a weak solution $(\mathfrak{F},X,W)$ to the SDE (1.1) with initial distribution $\mu$ in the sense of Definition 1.1, which is also unique in the class such that the following Krylov estimate holds: for any $(\alpha,p,q)\in[\alpha_{b},0]\times[2,\infty]^{2}$ with
    $\tfrac{q_{b}}{2}\leqslant q\leqslant q_{b},\quad p\leqslant p_{b},\quad\alpha-\tfrac{d}{p}-\tfrac{2}{q}\geqslant\alpha_{b}-\tfrac{d}{p_{b}}-\tfrac{2}{q_{b}},$ (1.6)
    and for any $T>0$, there exists $\theta>0$ such that for any $f\in{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}\cap{\mathbb{L}}^{q}_{T}C^{\infty}_{b}$ and $0\leqslant t_{0}<t_{1}\leqslant T$,
    $\left\|\int^{t_{1}}_{t_{0}}f(s,X_{s}){\mathord{{\rm d}}}s\right\|_{L^{m}(\Omega)}\leqslant C(t_{1}-t_{0})^{\frac{1+\theta}{2}}\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}},$ (1.7)
    where the constant $C$ is independent of $f$ and $t_{0},t_{1}$.
  2. (ii)
    If $c_{1}=0$, then the moment assumption on $\mu$ can be dropped; and for each $t>0$ and $X_{0}=x\in{\mathbb{R}}^{d}$, the law of $X_{t}$ admits a density $\rho_{t}(x,x^{\prime})$ called the heat kernel of $\Delta+b\cdot\nabla$, which satisfies the following two-sided Gaussian estimate: for fixed $T>0$, all $x,x^{\prime}\in{\mathbb{R}}^{d}$, and $t\in(0,T]$,
    $C_{0}t^{-d/2}\mathrm{e}^{-\gamma_{0}|x-x^{\prime}|^{2}/t}\leqslant\rho_{t}(x,x^{\prime})\leqslant C_{1}t^{-d/2}\mathrm{e}^{-\gamma_{1}|x-x^{\prime}|^{2}/t},$
    where $C_{0},C_{1},\gamma_{0},\gamma_{1}>0$ depend only on the parameters $T,d,\alpha_{b},p_{b},q_{b},c_{0}$, and $\|b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}}$.
@@
Under the linear growth assumption of $b_{2}$, we require a moment assumption on the initial distribution $\mu$, which arises from the Krylov estimates and the Young integral (see Proposition 4.4 and Lemma 4.6 below). Moreover, by Krylov’s estimate (1.7), one easily sees that
[…]
where the process $A^{b_{1}}$ defined in Definition 1.1 does not depend on the choice of the mollifiers.
Remark 1.4.
In the subcritical case, our result improves upon [19, Theorem 2.10] since our initial distribution is not required to have an $L^{2}$-density; it can be a Dirac measure. Moreover, we are working in the whole space, not in the torus. Our proof is based on the Schauder estimate of the heat equation in Besov spaces and Zvonkin’s transformation.
@@
(ii) If $b,\mathord{{\rm div}}b\in{\mathbf{B}}^{-1}_{\infty,2}$, then $b\in{\mathcal{B}}$. We shall prove it after introducing Besov spaces (see Lemma 2.7 below).
In the following we fix a time level $T>0$. Let ${\mathbb{C}}_{T}:=C([0,T];{\mathbb{R}}^{d})$ be the Banach space of all continuous functions. A path in ${\mathbb{C}}_{T}$ is denoted by $\omega$. The canonical process is denoted by $w_{t}(\omega)=\omega(t).$ Let ${\mathscr{B}}_{t}:=\sigma(w_{s},s\leqslant t)$ be the natural filtration.
@@
Let $\alpha_{b}\in[-1,0]$ and $p_{b},q_{b}\in[2,\infty]$ satisfy $\frac{2}{q_{b}}+\frac{d}{p_{b}}<2+\alpha_{b}$. Suppose that $b\in{\mathbb{L}}^{q_{b}}_{T}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}$ is divergence-free. For any initial distribution $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$ having $L^{2}$-density $\rho_{0}$, there is a weak solution $(\mathfrak{F},X,W)$ to SDE (1.1) in the sense of Definition 1.1, which has a density $\rho_{t}\in{\mathbb{L}}^{\infty}_{T}L^{2}\cap{\mathbb{L}}^{2}_{T}{\mathbf{H}}^{1}_{2}$ that satisfies the Fokker-Planck equation $\partial_{t}\rho=\Delta\rho-\mathord{{\rm div}}(b\rho)$ in the distributional sense. Moreover, consider the approximation SDE
@@
    ${\mathbb{E}}^{{\mathbb{P}}_{\mu}}[f(w_{t})|{\mathscr{B}}_{s}]={\mathbb{E}}^{{\mathbb{P}}_{\mu}}[f(w_{t})|w_{s}],\ \forall t\in[s,T],\ f\in C_{b}({\mathbb{R}}^{d}).$
@@
    If in addition that $b=b_{1}+b_{2}$, where $b_{1}\in{\mathbb{L}}^{\infty}_{T}{\mathcal{B}}$ and $b_{2}\in{\mathbb{L}}^{2}_{T}L^{2}$, then without subtracting subsequence, ${\mathbb{P}}_{n}$ weakly converges to ${\mathbb{P}}_{\mu}$ in ${\mathbb{C}}_{T}$ as $n\to\infty$, where ${\mathbb{P}}_{\mu}$ does not depend on the choice of the mollifiers.
@@
In the supercritical case, we are unable to demonstrate the uniqueness of weak solutions a priori. However, the assertion (A) affirms that we can identify a subsequence such that, for any initial value, the corresponding approximation solution converges weakly to a solution. This enables us to select a Markov process. In the case where $b=b_{1}+b_{2}$, with $b_{1}\in{\mathbb{L}}^{\infty}_{T}{\mathcal{B}}$ and $b_{2}\in{\mathbb{L}}^{2}_{T}L^{2}$, it is unique in the approximation sense since it is not necessary to choose a subsequence.
@@
When $f=0$ and $b=\mathord{{\rm div}}a$, where $a:\mathbb{R}^{d}\to\mathbb{R}^{d}\otimes\mathbb{R}^{d}$ is an antisymmetric matrix-valued measurable function, the Harnack inequality was established in [34] for $a\in L^{\infty}$ and in [37] for $a\in BMO$. In the case where $u_{0}=0$ and $b\in L^{p}$ with $\mathord{{\rm div}}b=0$, through De-Giorgi’s argument, the maximum estimate of weak solution $u$ was obtained in [51] for $p>d/2$ and in [48, Corollary 1.5] for $p>(d-1)/2$. In this work, for drifts that are distribution-valued and under certain conditions on the divergence of $b$, we show the weak well-posedness of PDE (1.9) using the energy method. To establish the existence of a weak solution to SDE (1.1), we rely on two types of Krylov’s estimates. One is used to prove tightness (see Lemma 5.1), while the other is utilized to take limits (see Lemma 5.5).
To establish conclusions (A) and (C), we will employ the concept of generalized martingale solutions. In scenarios involving distributional drifts, to the best of our knowledge, no existing results address the equivalence between weak solutions and generalized martingale solutions for SDEs with scaling supercritical distributional drifts. Nevertheless, in our proofs, both types of solutions are obtained as subsequential limits of smooth approximation solutions. It is worth emphasizing that in the supercritical case, even in the absence of uniqueness, we retain the flexibility to select a Markov process.
@@
In [25, Proposition 1], the entropy solution of the above Liouville equation is directly constructed. However, the existence of a solution to the SDE (1.10) appears to be an open question. In the context of the two-dimensional vortex model, where